a bookshelf is 36 inches long. It has n book on it. If each book were reduced by 1'2 inch, an additional 6 books could be held on the shelf =w.
What is N?.
What is W.?
Thank you

Answer by checkley75(3666)

36/W=N (WIDTH OF THE BOOKS)
36/(W-.5)=N+6 SUBSTITUTE (36/W) FOR N & SOLVE FOR W
36/(W-.5)=36/W+6
36/(W-.5)=(36+6W)/W NOW CROSS MULTIPLY
36W=(W-.5)(36+6W)
36W=36W-18+6W^2-3W
6W^2-3W-18=0
2W^2-W-6=0
(2W+3)(W-2)=0
2W+3=0
2W=-3
W=-3/2
W-2=0
W=2 ANSWER. THE WIDTH OF THE BOOKS. THUS:
36/2=18 BOOKS FIT THE SHELF.
36/1.5=24 WHICH IS 6 MORE BOOKS IF THE BOOKS ARE REDUCED BY .5 INCHES.